OpenStudy (anonymous):

what number should be added to x^2 - 4x to complete the square????????????

jhonyy9 (jhonyy9):

how do you think it ?

OpenStudy (anonymous):

Note (a+b)^2 = a^2+2ab+b^2; the middle term 2ab corresponds to our -4x. Since a is x, we see b must be -2. To determine the third term we add the square of b, i.e. (-2)^2 = 4, so x^2 - 4x + 4 = (x - 2)^2

OpenStudy (anonymous):

try to figur it out. still trying. plugging stuff in.

OpenStudy (anonymous):

trying

OpenStudy (anonymous):

is it 2

OpenStudy (anonymous):

Did you read my post? One must add 4...

OpenStudy (anonymous):

6

OpenStudy (anonymous):

The answer is neither 2 nor 6. I've stated it twice.

OpenStudy (anonymous):

oh its 4

OpenStudy (anonymous):

Think of it like this... (a+b)^2 = a^2 + 2ab + b^2 We're given the fragment x^2 - 4x; these two terms correspond to two of a^2 + 2ab + b^2. Since a^2 = x^2, therefore a = x; similarly, we know 2ab = -4x, i.e. 2b = -4. This leaves b = -2. Now, we need to add that third term -- b^2 -- to complete the square, and b^2 = (-2)^2 = 4. x^2 - 4x + 4 = (x - 2)^2